OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 36 inches and a height of 24 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (xapproachesinfinity):

the side length and height were increased by the factor of 3 so you would have volume also increased by a factor of 3^3

OpenStudy (xapproachesinfinity):

Volume of a pyramid \(V=B^2H/3\) B=3b H=3h where B is the base length of the large pyramid and H is its height so \(V=9b^23h/3=9b^2h=(3^2)b^2h\)